OpenStudy (anonymous):

Use the distributive property to simplify the expression below: 3x-21/3

mathslover (mathslover):

is the question : \[\huge{3(3x-21)}\] ?

mathslover (mathslover):

or \[\huge{\frac{3x-21}{3}}\]

mathslover (mathslover):

OK ! so here i go ! FIRST OF ALL : \[\huge{\textbf{Welcome to openstudy}}\] assuming that this is ur question : \[\huge{\frac{3x-21}{3}}\] \[\huge{\frac{3(x)-(7)(3)}{3}}\] \[\huge{\frac{3(x-7)}{3}}\] Can u simplify this further ? @zahabaha

OpenStudy (anonymous):

Choose the Standard Form of the equation y = −x − 6

mathslover (mathslover):

i request u to post this as a new question ..

OpenStudy (anonymous):

y=-14/x-6 plz answer